# Inversion-DPO: Precise and Efficient Post-Training for Diffusion Models

### v2

@@ 3.2 DDIM Inversion
Figure 1: Computation pipeline of the proposed Inversion-DPO. 1) Given an initial image $x_{0}$, we obtain the diffusion trajectories $\{x_{1},x_{2},\ldots,x_{T}\}$ of both the winning and losing samples from a reference model via DDIM inversion. 2) At each timestep along the diffusion trajectory, the model is trained to predict the noise $\epsilon$, aligning it closely with the $\epsilon$ of the winning sample while diverging from that of the losing sample. 3) This process is repeated across the entire diffusion trajectory, continuously updating model parameters until the final predicted image is generated.
4.1 Inversion Direct Preference Optimization
@@
DPO [45] is an effective reward-guided post-training approach that directly learns human preferences through a reparameterized policy, rather than optimizing a reward function and then performing reinforcement learning (RL).
@@
Recall that when adapting DPO to diffusion models [57] , the objective shifts from fitting $x_{0}$ to fitting the trajectory $(x_{0},\dots,x_{T})$with
@@
$(\displaystyle\mathcal{L}_{\text{Diffusion-DPO}}(\theta)=-\mathbb{E}_{(\boldsymbol{x}_{0}^{w},\boldsymbol{x}_{0}^{l})}\log\sigma\bigg{(}$10)
@@
We address these issues by designing Inversion-DPO (Fig. 1). Our key insight stems from the deterministic characteristic of DDIM Inversion [52]. Given a sample $\boldsymbol{x}_{0}$, DDIM Inversion constructs the sampling trajectory $\boldsymbol{x}_{1:T}$ leading to $\boldsymbol{x}_{0}$ through a deterministic mapping by equation (9). Thus, we propose to approximate the posterior $p_{\theta}(\boldsymbol{x}_{1:T}|\boldsymbol{x}_{0})$ with DDIM inversion. This eliminates the need for stochastic approximation $q$ in trajectory estimation. Using DDIM Inversion is reasonable with three advantages. First, assuming the deterministic sampling setting, the posterior inference is the inversion process. Thus, the resulting $\boldsymbol{x}_{1:T}$ are more precise. Second, such inversion is compatible with the definition of the reward function on diffusion in equation (5), so the implicit reward learning is also precise. Third, the final estimated expectation in (10) is also more precise in this case.
@@
$\log\frac{p_{\theta}(\boldsymbol{x}_{0:T}^{w})}{p_{\theta_{0}}(\boldsymbol{x}_{0:T}^{w})}=\sum_{t=1}^{T}\left[\log p_{\theta}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})-\log p_{\theta_{0}}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})\right]$ (11)
Since $p_{\theta}$ and $p_{\theta_{0}}$ share the same distribution of $x_{T}$ as standard Gaussian, $\log p_{\theta}(x_{T}^{w})-\log p_{\theta_{0}}(x_{T}^{w})$ cancel out. The conditional probabilities of $p_{\theta}$ and $p_{\theta_{0}}$ are Gaussian distributions as in equation (8) and becomes dirac-delta as $\sigma_{t}$ approaches 00. We first continue derivation with $\sigma_{t}\neq 0$ and adopt a simple approximation [17] of the final loss.
@@
$\boldsymbol{\mu}_{\theta_{0}}(x_{t}^{w})=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}\boldsymbol{x}_{t}^{w}+$ (12)
$\left(\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}-\frac{\sqrt{\alpha_{t-1}({1-\alpha_{t}})}}{\sqrt{\alpha_{t}}}\right)\epsilon_{\theta_{0}}(\boldsymbol{x}_{t}^{w},t)$ (13)
@@
$-\log p_{\theta_{0}}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})$ $\propto\|\boldsymbol{\mu}_{\theta_{0}}(x_{t}^{w})-\boldsymbol{x}_{t-1}^{w}\|^{2}$ (14)
@@
$\log\frac{p_{\theta}(\boldsymbol{x}_{0:T}^{w})}{p_{\theta_{0}}(\boldsymbol{x}_{0:T}^{w})}\propto\left[\sum_{t=1}^{T}\|\epsilon_{\theta}(\boldsymbol{x}_{t}^{w},t)-\epsilon_{\theta_{0}}(\boldsymbol{x}_{t}^{w},t)\|^{2}\right],$ (15)
@@
$\mathcal{L}_{\text{Inversion-DPO}}(\theta)=-\mathbb{E}_{(\boldsymbol{x}_{0}^{w},\boldsymbol{x}_{0}^{l})}\log\sigma\bigg{(}\beta\mathbb{E}_{\begin{subarray}{c}\boldsymbol{x}_{1:T}^{w}\sim p_{\theta}(\boldsymbol{x}_{1:T}^{w}|\boldsymbol{x}_{0}^{w})\\ \boldsymbol{x}_{1:T}^{l}\sim p_{\theta}(\boldsymbol{x}_{1:T}^{l}|\boldsymbol{x}_{0}^{l})\end{subarray}}$ (16)
@@
Notably, during the training of Inversion-DPO, the previous four-term loss collapses into two deterministic summations due to trajectory determinism, which is conceptually simpler. This simplification also yields non-negligible improvements in optimization efficiency for model training. The pretrained $\epsilon_{\theta_{0}}$ can be transformed to approximate the score. Fitting the noise added as in (3.1) introduces extra variance [56, 41] which in unstable in the training [64], while fitting the predicted noise close to the score is more stable and thus improves the training efficiency. Our training schema fundamentally resolves the precision-efficiency tradeoff in preference-based diffusion alignment.
4.2 Inversion-DPO with multiple objectives
For effective Inversion-DPO training, both positive (i.e., winning samples) and negative (i.e., losing samples) data pairs are required to guide the optimization. While manual annotation is commonly used to obtain such pairs [71, 57], it is costly and prone to ambiguity. To address this, recent works [54, 3, 65, 39] adopt automatic metrics to rank images as guidance. We follow this strategy. However, relying on a single metric is often inadequate to comprehensively capture the quality of the outputs, especially in complex scene generation scenarios. Therefore, we introduce a multi-objective learning strategy, leveraging multiple metrics to provide more reliable and informative preference signals.
@@ 4.2 Inversion-DPO with multiple objectives
$r(c,x_{0})=\frac{1}{N}\sum_{i=1}^{N}r_{i}\left(c,x_{0}\right)$ (17)
@@
We evaluate the effectiveness of Inversion-DPO in two representative tasks. In addition to the typical text-to-image generation, we further assess it in the more challenging setting of compositional image generation, which requires the model to accurately synthesize complex scenes involving multiple objects and relationships. Existing studies [31, 59] suggest that, compared to textual annotations, structured annotations provide a more compact and explicit form of information integration, making them more effective for representing complex scenes. Such structured annotations are typically represented as graphs, consisting of nodes and edges that denote objects in the scene and their interactions, respectively.
@@
On basic text-to-image tasks, Pick-a-Pic dataset [26] serves as a reliable source of supervision for training Inversion-DPO. However, for compositional generation, there is a lack of corresponding paired datasets that can be directly utilized. This is because existing datasets with structured annotations [4, 27, 31] provide only single-condition samples, making them insufficient for preference-based learning. To bridge this gap, we construct a paired dataset with structured annotations to support alignment and evaluation.
@@
In detail, for a given complex scene annotation, we generate four samples using an existing baseline [31], and compute a multi-dimensional score for each generated image. During training, the model determines positive and negative samples based on a combination of these scores, guiding the learning of Inversion-DPO.
@@
To ensure effective evaluation of complex scene images, we adopt two representative metrics–DPG-Bench Score [18] and the 3-in-1 score from T2I-CompBench [19]–as the criteria for multi-objective learning. In total, we generate 11,140 images for 2,785 annotations, with each image associated with a comprehensive complex scene score. Rather than enforcing fixed pairwise bindings, we allow the model to dynamically determine relative “winning” and “losing” samples based on their comprehensive scores during training. This approach prevents the dataset size from being constrained to the original 2,785 samples, enabling scalable data expansion within a certain range and ultimately improving training quality.
5 Experiments
@@
Figure 2: A demonstration of how PickScore increase with more training steps. The blue curve shows the performance of our Inversion-DPO trained with different numbers of DDIM inversion steps. The gray curve represents our retraining with the official implementation of Diffusion-DPO [57], while the dashed line corresponds to the officially released weights. Our model with 80-step inversion outperforms the both the retraining and the released Diffusion-DPO checkpoint. All models are trained under identical configurations.
5.1 Dataset and Implementation Details
@@
Dataset. We use the Pick-a-Pic v2 dataset [26] to train our T2I model, which contains paired preferences of images generated by SDXL-beta and Dreamlike, along with prompts and preference pairs collected from the web. Following the same data processing approach as Diffusion-DPO [57], we exclud 12% of tie samples, ultimately obtaining 851,293 effective data pairs. For the compositional image generation task, as described in Section 4.2, we curate compositional image pairs from LAION-SG [31] for training, ensuring data quality and diversity.
@@
Training. During the training process, we employ the Adafactor [50] optimizer for SDXL and SGXL-SG to save memory, while AdamW [36] was used for SD1.5 training. The learning rate for all models was uniformly set to 1e-6. Specifically, the $\beta$ parameter in Inversion-DPO was set to 2000. We also set the Inversion Steps to 20, 40, and 80, respectively, to observe the impact of different DDIM Inversion steps on model performance, which is detailed in our ablation study. Each model was trained for 2000 steps, with all training tasks completed on two NVIDIA H800 GPUs.
5.2 Baselines and Evaluation Metrics
@@
For the typical text-to-image generation task, we evaluate a set of representative baseline models and post-training methods, including SD1.5 [47], SDXL [43], Diffusion-DPO [57], DDPO [3], D3PO [65], Demon [69], and IterComp [71]. The generated results are assessed using PickScore [26], CLIP Score [44], and Aesthetic Score [49], which respectively measure alignment with human preference, semantic consistency with the input prompt, and the aesthetic quality of image details.
@@
For compositional image generation, we evaluate SGDiff [66], SG-Adapter [51], SDXL-SG [31], and SD1.5-SG [31]. Considering the objectives of complex scene generation, we adopt FID [16] as the image quality metric, along with SG-IoU, Entity-IoU, and Relation-IoU [51], which respectively assess scene-level consistency, object-level consistency, and the consistency of relationships between objects within the generated images.
5.3 Qualitative Results
@@
Figure 3: A comparison of generation details. Existing models often struggle to produce fine-grained details, while Inversion-DPO demonstrates a higher level of fidelity in generating nuanced elements such as limbs, ties, and textures.
Figure 4: Qualitative results on compositional image generation. Inversion-DPO proves effective in addressing key challenges in compositional image generation, including object omission (orange and green annotations in (a), (d), and (e)), attribute errors (red annotations in (b), (c), and (d)), and relation inaccuracies (blue annotations in (c) and (d)). As shown in the last row, Inversion-DPO achieves satisfactory generation results on cases where baseline models struggle.
@@
LABEL:fig:_teaser illustrates the visual performance of our Inversion-DPO compared to baseline models on text-to-image tasks. Each row shows an input prompt and the corresponding generated images. Compared to existing baselines [43, 3, 69, 57], Inversion-DPO exhibits higher visual appeal and achieves higher PickScore. Moreover, in Fig. 3, our method demonstrates promising potential in controlling fine-grained detail generation, which is a persistent challenge for text-to-image models. We further visualize results on more challenging compositional image generation tasks. As shown in Fig. 4, Inversion-DPO demonstrates superior performance in terms of object occurrence, relation accuracy, and attribute binding.

### v4

@@ 4. Method
Figure 2. Computation pipeline of the proposed Inversion-DPO. 1) Given an initial image $x_{0}$, we obtain the diffusion trajectories $\{x_{1},x_{2},\ldots,x_{T}\}$ of both the winning and losing samples from a reference model via DDIM inversion. 2) At each timestep along the diffusion trajectory, the model is trained to predict the noise $\epsilon$, aligning it closely with the $\epsilon$ of the winning sample while diverging from that of the losing sample. 3) This process is repeated across the entire diffusion trajectory, continuously updating model parameters until the final predicted image is generated.
4.1. Inversion Direct Preference Optimization
@@
DPO (Rafailov et al., 2023) is an effective reward-guided post-training approach that directly learns human preferences through a reparameterized policy, rather than optimizing a reward function and then performing reinforcement learning (RL).
@@
Recall that when adapting DPO to diffusion models (Wallace et al., 2024) , the objective shifts from fitting $x_{0}$ to fitting the trajectory $(x_{0},\dots,x_{T})$with
@@
(10) $\mathcal{L}_{\text{Diffusion-DPO}}(\theta)=-\mathbb{E}_{(\boldsymbol{x}_{0}^{w},\boldsymbol{x}_{0}^{l})}\log\sigma\bigg{(}$
@@
We address these issues by designing Inversion-DPO (fig. 2). Our key insight stems from the deterministic characteristic of DDIM Inversion (Song et al., 2020). Given a sample $\boldsymbol{x}_{0}$, DDIM Inversion constructs the sampling trajectory $\boldsymbol{x}_{1:T}$ leading to $\boldsymbol{x}_{0}$ through a deterministic mapping by equation (9). Thus, we propose to approximate the posterior $p_{\theta}(\boldsymbol{x}_{1:T}|\boldsymbol{x}_{0})$ with DDIM inversion. This eliminates the need for stochastic approximation $q$ in trajectory estimation. Using DDIM Inversion is reasonable with three advantages. First, assuming the deterministic sampling setting, the posterior inference is the inversion process. Thus, the resulting $\boldsymbol{x}_{1:T}$ are more precise. Second, such inversion is compatible with the definition of the reward function on diffusion in equation (5), so the implicit reward learning is also precise. Third, the final estimated expectation in (10) is also more precise in this case.
@@
(11) $\log\frac{p_{\theta}(\boldsymbol{x}_{0:T}^{w})}{p_{\theta_{0}}(\boldsymbol{x}_{0:T}^{w})}=\sum_{t=1}^{T}\left[\log p_{\theta}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})-\log p_{\theta_{0}}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})\right]$
Since $p_{\theta}$ and $p_{\theta_{0}}$ share the same distribution of $x_{T}$ as standard Gaussian, $\log p_{\theta}(x_{T}^{w})-\log p_{\theta_{0}}(x_{T}^{w})$ cancel out. The conditional probabilities of $p_{\theta}$ and $p_{\theta_{0}}$ are Gaussian distributions as in equation (8) and becomes dirac-delta as $\sigma_{t}$ approaches 0. We first continue derivation with $\sigma_{t}\neq 0$ and adopt a simple approximation (Ho et al., 2020) of the final loss.
@@
$\boldsymbol{\mu}_{\theta_{0}}(x_{t}^{w})=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}\boldsymbol{x}_{t}^{w}+$
(12) $\left(\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}-\frac{\sqrt{\alpha_{t-1}({1-\alpha_{t}})}}{\sqrt{\alpha_{t}}}\right)\epsilon_{\theta_{0}}(\boldsymbol{x}_{t}^{w},t)$
@@
(13) $-\log p_{\theta_{0}}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})$ $\propto\|\boldsymbol{\mu}_{\theta_{0}}(x_{t}^{w})-\boldsymbol{x}_{t-1}^{w}\|^{2}$
@@
(14) $\log\frac{p_{\theta}(\boldsymbol{x}_{0:T}^{w})}{p_{\theta_{0}}(\boldsymbol{x}_{0:T}^{w})}\propto\left[\sum_{t=1}^{T}\|\epsilon_{\theta}(\boldsymbol{x}_{t}^{w},t)-\epsilon_{\theta_{0}}(\boldsymbol{x}_{t}^{w},t)\|^{2}\right],$
@@
(15) $\mathcal{L}_{\text{Inversion-DPO}}(\theta)=-\mathbb{E}_{(\boldsymbol{x}_{0}^{w},\boldsymbol{x}_{0}^{l})}\log\sigma\bigg{(}\beta\mathbb{E}_{\begin{subarray}{c}\boldsymbol{x}_{1:T}^{w}\sim p_{\theta}(\boldsymbol{x}_{1:T}^{w}|\boldsymbol{x}_{0}^{w})\\ \boldsymbol{x}_{1:T}^{l}\sim p_{\theta}(\boldsymbol{x}_{1:T}^{l}|\boldsymbol{x}_{0}^{l})\end{subarray}}$
@@
Notably, during the training of Inversion-DPO, the previous four-term loss collapses into two deterministic summations due to trajectory determinism, which is conceptually simpler. This simplification also yields non-negligible improvements in optimization efficiency for model training. The pretrained $\epsilon_{\theta_{0}}$ can be transformed to approximate the score. Fitting the noise added as in (3.1) introduces extra variance (Vincent, 2011; Mittal et al., 2023) which in unstable in the training (Xu et al., 2023), while fitting the predicted noise close to the score is more stable and thus improves the training efficiency. Our training schema fundamentally resolves the precision-efficiency tradeoff in preference-based diffusion alignment.
4.2. Inversion-DPO with Multiple Objectives
For effective Inversion-DPO training, both positive (i.e., winning samples) and negative (i.e., losing samples) data pairs are required to guide the optimization. While manual annotation is commonly used to obtain such pairs (Zhang et al., 2024c; Wallace et al., 2024), it is costly and prone to ambiguity. To address this, recent works (Tian et al., 2025; Black et al., 2024; Yang et al., 2024b; Ma et al., 2025) adopt automatic metrics to rank images as guidance. We follow this strategy. However, relying on a single metric is often inadequate to comprehensively capture the quality of the outputs, especially in complex scene generation scenarios. Therefore, we introduce a multi-objective learning strategy, leveraging multiple metrics to provide more reliable and informative preference signals.
@@ 4.2. Inversion-DPO with Multiple Objectives
(16) $r(c,x_{0})=\frac{1}{N}\sum_{i=1}^{N}r_{i}\left(c,x_{0}\right)$
@@
We evaluate the effectiveness of Inversion-DPO in two representative tasks. In addition to the typical text-to-image generation, we further assess it in the more challenging setting of compositional image generation, which requires the model to accurately synthesize complex scenes involving multiple objects and relationships. Existing studies (Li et al., 2024b; Wang et al., 2024b) suggest that, compared to textual annotations, structured annotations provide a more compact and explicit form of information integration, making them more effective for representing complex scenes. Such structured annotations are typically represented as graphs, consisting of nodes and edges that denote objects in the scene and their interactions, respectively.
@@
On basic text-to-image tasks, Pick-a-Pic dataset (Kirstain et al., 2023) serves as a reliable source of supervision for training Inversion-DPO. However, for compositional generation, there is a lack of corresponding paired datasets that can be directly utilized. This is because existing datasets with structured annotations (Caesar et al., 2018; Krishna et al., 2017; Li et al., 2024b) provide only single-condition samples, making them insufficient for preference-based learning. To bridge this gap, we construct a paired dataset with structured annotations to support alignment and evaluation.
@@
In detail, for a given complex scene annotation, we generate four samples using an existing baseline (Li et al., 2024b), and compute a multi-dimensional score for each generated image. During training, the model determines positive and negative samples based on a combination of these scores, guiding the learning of Inversion-DPO.
@@
To ensure effective evaluation of complex scene images, we adopt two representative metrics–DPG-Bench Score (Hu et al., 2024) and the 3-in-1 score from T2I-CompBench (Huang et al., 2023)–as the criteria for multi-objective learning. In total, we generate 11,140 images for 2,785 annotations, with each image associated with a comprehensive complex scene score. Rather than enforcing fixed pairwise bindings, we allow the model to dynamically determine relative “winning” and “losing” samples based on their comprehensive scores during training. This approach prevents the dataset size from being constrained to the original 2,785 samples, enabling scalable data expansion within a certain range and ultimately improving training quality.
5. Experiments
@@
Figure 3. A demonstration of how PickScore increase with more training steps. The blue curve shows the performance of our Inversion-DPO trained with different numbers of DDIM inversion steps. The gray curve represents our retraining with the official implementation of Diffusion-DPO (Wallace et al., 2024), while the dashed line corresponds to the officially released weights. Our model with 80-step inversion outperforms the both the retraining and the released Diffusion-DPO checkpoint. All models are trained under identical configurations.
Figure 4. Relationship between training steps and Aesthetics Score improvements. The orange curve shows the performance of our Inversion-DPO trained with different numbers of DDIM inversion steps. The gray curve represents our implementation of Diffusion-DPO (Wallace et al., 2024), while the dashed line corresponds to the officially released weights of Diffusion-DPO. All models are trained under identical configurations.
5.1. Dataset and Implementation Details
@@
Dataset. We use the Pick-a-Pic v2 dataset (Kirstain et al., 2023) to train our T2I model, which contains paired preferences of images generated by SDXL-beta and Dreamlike, along with prompts and preference pairs collected from the web. Following the same data processing approach as Diffusion-DPO (Wallace et al., 2024), we exclud 12% of tie samples, ultimately obtaining 851,293 effective data pairs. For the compositional image generation task, as described in Section 4.2, we curate compositional image pairs from LAION-SG (Li et al., 2024b) for training, ensuring data quality and diversity.
@@
Training. During the training process, we employ the Adafactor (Shazeer and Stern, 2018) optimizer for SDXL and SGXL-SG to save memory, while AdamW (Loshchilov and Hutter, 2019) was used for SD1.5 training. The learning rate for all models was uniformly set to 1e-6. Specifically, the $\beta$ parameter in Inversion-DPO was set to 2000. We also set the Inversion Steps to 20, 40, and 80, respectively, to observe the impact of different DDIM Inversion steps on model performance, which is detailed in our ablation study. Each model was trained for 2000 steps, with all training tasks completed on two NVIDIA H800 GPUs.
5.2. Baselines and Evaluation Metrics
@@
For the typical text-to-image generation task, we evaluate a set of representative baseline models and post-training methods, including SD1.5 (Rombach et al., 2022), SDXL (Podell et al., 2023), Diffusion-DPO (Wallace et al., 2024), DDPO (Black et al., 2024), D3PO (Yang et al., 2024b), Demon (Yeh et al., 2025), and IterComp (Zhang et al., 2024c). The generated results are assessed using PickScore (Kirstain et al., 2023), CLIP Score (Radford et al., 2021), and Aesthetic Score (Schuhmann et al., 2022), which respectively measure alignment with human preference, semantic consistency with the input prompt, and the aesthetic quality of image details.
@@
For compositional image generation, we evaluate SGDiff (Yang et al., 2022), SG-Adapter (Shen et al., 2024), SDXL-SG (Li et al., 2024b), and SD1.5-SG (Li et al., 2024b). Considering the objectives of complex scene generation, we adopt FID (Heusel et al., 2017) as the image quality metric, along with SG-IoU, Entity-IoU, and Relation-IoU (Shen et al., 2024), which respectively assess scene-level consistency, object-level consistency, and the consistency of relationships between objects within the generated images.
5.3. Qualitative Results
@@
Figure 5. A comparison of generation details. Existing models often struggle to produce fine-grained details, while Inversion-DPO demonstrates a higher level of fidelity in generating nuanced elements such as limbs, ties, and textures.
@@
Figure 1 illustrates the visual performance of our Inversion-DPO compared to baseline models on text-to-image tasks. Each row shows an input prompt and the corresponding generated images. Compared to existing baselines (Podell et al., 2023; Black et al., 2024; Yeh et al., 2025; Wallace et al., 2024), Inversion-DPO exhibits higher visual appeal and achieves higher PickScore. Moreover, in fig. 5, our method demonstrates promising potential in controlling fine-grained detail generation, which is a persistent challenge for text-to-image models. We further visualize results on more challenging compositional image generation tasks. Moreover, Inversion-DPO demonstrates superior performance in terms of object occurrence, relation accuracy, and attribute binding (details included in supplementary material).